# Classification of abelian Schur groups I

### v1

@@ 7.5. Auxiliary statements
Due to Lemma 6.2 and Lemma 8.2, the $S$-ring $\mathcal{A}_{S_{D}}$ is normal or trivial over a group of order $p$. If $\mathcal{A}_{S_{D}}$ is normal or $|S_{D}|=3$, then $\mathcal{A}_{S_{D}}$ is $2$-minimal by Lemma 6.5 or Lemma 2.6, respectively. So $K_{1}^{S_{D}}=K_{0}^{S_{D}}=\operatorname{Aut}(\mathcal{A}_{S_{D}})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ as required. Further, we assume that
@@
By Lemma 8.5, we have $K_{1}^{S_{H}}=K_{0}^{S_{H}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}$ is not the $T$-wreath product, where $T=U_{1}/L_{1}\in\{(H\times L)/L,U/(U\cap D)\}$, then Lemma 8.6 implies that $K_{1}^{S_{D}}=K_{0}^{S_{D}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$. Therefore $K_{1}^{S}=K_{0}^{S}$ by Lemma 3.3. Thus, $\mathcal{A}$ is schurian by Lemma 3.5.
Suppose that $\mathcal{A}$ is the $T$-wreath product. Clearly, $|T|=|H|=4$ and Statement $(4)$ of Theorem 7.1 holds for $\mathcal{A}$ and $T$. Lemma 8.5 yields that there are $M_{1}\leq\operatorname{Aut}(\mathcal{A}_{U_{1}})$ and $M_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$ such that $M_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U_{1}})$, $M_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$, and $M_{1}^{T}=M_{0}^{T}$. Again, $\mathcal{A}$ is schurian by Lemma 3.5.
@@
Since a section of a Schur group is a Schur group, to prove Theorem 1.3, it suffices to prove that the group $C_{8}\times C_{2}\times C_{p}$, where $p$ is an odd prime, is not a Schur group. Let $A\cong C_{8}$, $B\cong C_{2}$, $P\cong C_{p}$, and $G=A\times B\times P\cong C_{8}\times C_{2}\times C_{p}$. Denote generators of $A$ and $B$ by $a$ and $b$, respectively. Put $A_{1}=\langle a^{2}\rangle\cong C_{4}$, $A_{2}=\langle a^{4}\rangle\cong C_{2}$, $U_{1}=A_{1}\times P\cong C_{4p}$, and $U=A_{1}\times B\times P=B\times U_{1}\cong C_{2}\times C_{4p}$.
Let us construct a nonschurian $S$-ring over $G$. The group $M=\operatorname{Aut}(P)$ is a cyclic group of order p1p-1. Let $M_{0}$ be the subgroup of $M$ of index $2$. The canonical epimorphism from $M$ to $M/M_{0}$ is denoted by $\pi$. The group $N=\operatorname{Aut}(A_{1})$ is a cyclic group of order $2$. There exists a unique isomorphism $\varphi$ from $M/M_{0}\cong C_{2}$ to $N\cong C_{2}$. Put
$F=\{(\sigma,\tau)\in M\times N:~(\sigma^{\pi})^{\varphi}=\tau\}\leq\operatorname{Aut}(U_{1})$
and $\mathcal{A}_{1}=\operatorname{Cyc}(F,U_{1})$. Let $\mathcal{A}_{2}$ be an $S$-ring over $B\times P$ such that $\mathcal{A}_{2}=\mathcal{T}_{B}\otimes\mathcal{T}_{P}$.
@@
Let $\mathcal{A}_{3}=\operatorname{Cyc}(\langle\sigma_{0}\rangle,A)$, where $\sigma_{0}\in A$ is such that $a^{\sigma_{0}}=a^{-1}$. It is easy to check in a straightforward way that the partition of $(A/A_{2})\times B\cong C_{4}\times C_{2}$ into the sets
@@
defines the $S$-ring $\mathcal{A}_{4}$ over $(A/A_{2})\times B$. In fact,
@@
Since $X_{3}=P^{\#}$ is an orbit of $K$, we conclude that
@@
Let $a,b,c,d\in H$ be such that $\langle a,b,c,d\rangle=H$, $A=\langle a\rangle$, $B=\langle b\rangle$, $C=\langle c\rangle$, and $\sigma_{0}\in\operatorname{Aut}(H)$ such that
$\sigma_{0}:(a,b,c,d)\mapsto(a,ab,bc,cd),$
$M_{0}=\langle\sigma_{0}\rangle$, and $\mathcal{A}_{0}=\operatorname{Cyc}(M_{0},H)$. It is easy to verify in a straightforward way that $|\sigma_{0}|=4$ and the basic sets of $\mathcal{A}_{0}$ are the following:
@@
One can see that $\mathcal{A}_{0}$ is a 22-$S$-ring, i.e. a size of every basic set is a $2$-power, and $\mathcal{A}_{0}$ is indecomposable. Therefore $\mathcal{A}_{0}$ is $2$-minimal by [17, Lemma 5.6]. Thus,
@@
for every $K_{0}\leq\operatorname{Sym}(H)$ such that $K_{0}\geq H_{r}$ and $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0})$. In particular, $\operatorname{Aut}(\mathcal{A}_{0})=H_{r}\rtimes M_{0}$.
Let $V=A\times B\times C$ and $U=V\times P$. As $p\geq 5$, the group $M_{P}=\operatorname{Aut}(P)\cong C_{p-1}$ has a nontrivial subgroup $M_{P}^{0}$ of index $2$. Let $\sigma_{1},\sigma_{2}\in\operatorname{Aut}(V)$ be such that
$\sigma_{1}:(a,b,c)\mapsto(a,ab,bc),~\sigma_{2}:(a,b,c)\mapsto(a,b,bc).$
Put $M_{V}=\langle\sigma_{1},\sigma_{2}\rangle$. It can be verified straightforwardly that $|\sigma_{1}|=4$, $|\sigma_{2}|=2$, and $\sigma_{2}\sigma_{1}\sigma_{2}=\sigma_{1}^{-1}$. So $M_{V}\cong D_{8}$.
Let $M^{0}_{V}=\langle\sigma_{1}^{2},\sigma_{2}\sigma_{1}\rangle$. Clearly, $M_{V}^{0}\cong E_{4}$ and hence $M_{V}/M_{V}^{0}\cong C_{2}$. Put
$M_{1}=\{(\sigma,\tau)\in M_{V}\times M_{P}:~(\sigma^{\pi_{1}})^{\psi}=\tau^{\pi_{2}}\}\leq\operatorname{Aut}(U),$
where $\pi_{1}$ and $\pi_{2}$ are the canonical epimorphisms from $M_{V}$ to $M_{V}/M_{V}^{0}$ and from $M_{P}$ to $M_{P}/M_{P}^{0}$, respectively, and $\psi$ is a unique isomorphism from $M_{V}/M_{V}^{0}\cong C_{2}$ to $M_{P}/M_{P}^{0}\cong C_{2}$, and $\mathcal{A}_{1}=\operatorname{Cyc}(M_{1},U)$. Due to the definition of $M_{1}$, we conclude that $V$ and $P$ are $\mathcal{A}_{1}$-subgroups and the basic sets of $\mathcal{A}_{1}$ are the following:
@@
$K_{1}\geq U_{r},~K_{0}\geq(G/L)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{1}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0}),~\text{and}~K_{0}^{S}=K_{1}^{S}$
@@
$X=AcP_{1}\cup AbcP_{2},$
$x\in P_{1}$, and $f\in M_{x}$, where $M=(K_{1})_{e}$. Due to $K_{0}^{S}=K_{1}^{S}$ and Eq. (38), we have $f^{S}\in M_{0}^{S}$. Eq. (2) and inclusions $(A\times B)c,X^{-1}\in\mathcal{S}(\mathcal{A})$ imply that
@@
Together with $f^{S}\in M_{0}^{S}$ and the definition of $M_{0}$, this yields that $f^{S}\in\langle\sigma_{0}^{2}\rangle$ and hence $b^{f}=b$. Since $H$ and $P$ are $\mathcal{A}$-subgroups, $Hx$ and $Pb$ are blocks of $\operatorname{Aut}(\mathcal{A})$. As $x^{f}=x$ and $b^{f}=b$, we conclude that $(Hx)^{f}=Hx$ and $(Pb)^{f}=Pb$. Therefore
@@
Thus, $f\in M_{bx}$ and consequently
@@
The definition of $M_{1}$ implies that the basic set of $\mathcal{A}$ containing $x$ is $P^{\#}$ and the basic set of $\mathcal{A}$ containing $bx$ is $Y=AbP^{\#}$. Clearly, |Y|=2(p1)|Y|=2(p-1). Due to the assumption, we have $P^{\#},Y\in\operatorname{Orb}(M,U)$. Therefore
@@
a contradiction to Eq. (39). ∎
Now Theorem 1.4 follows from Proposition 10.1.
Acknowledgment: The author is very grateful to Dr. Matan Ziv-Av for the help with computer calculations.

### v3

@@ 7.5. Auxiliary statements
By Lemma 6.2 and Lemma 8.2, the $S$-ring $\mathcal{A}_{S_{D}}$ is normal or trivial over a group of order $p$. If $\mathcal{A}_{S_{D}}$ is normal, then $\mathcal{A}_{S_{D}}$ is $2$-minimal by Lemma 6.5. If $|S_{D}|=3$, then $\mathcal{A}_{S_{D}}$ is obviously $2$-minimal. So $K_{1}^{S_{D}}=K_{0}^{S_{D}}=\operatorname{Aut}(\mathcal{A}_{S_{D}})\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$ in both these cases as required. Further, we assume that
@@
Return to the proof of Theorem 1.2. By Lemma 8.5, we have $K_{1}^{S_{H}}=K_{0}^{S_{H}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{H}})$. If $\mathcal{A}$ is not the $T$-wreath product, where $T=U_{1}/L_{1}\in\{(H\times L)/L,U/(U\cap D)\}$, then Lemma 8.6 implies that $K_{1}^{S_{D}}=K_{0}^{S_{D}}\in\mathcal{K}^{\min}(\mathcal{A}_{S_{D}})$. Therefore $K_{1}^{S}=K_{0}^{S}$ by Lemma 3.3. Thus, $\mathcal{A}$ is schurian by Lemma 3.5.
Suppose that $\mathcal{A}$ is the $T$-wreath product. Clearly, $|T|=|H|=4$ and Statement $(4)$ of Theorem 7.1 holds for $\mathcal{A}$ and $T$. By Lemma 8.5, there are $M_{1}\leq\operatorname{Aut}(\mathcal{A}_{U_{1}})$ and $M_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$ such that $M_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{U_{1}})$, $M_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{G/L_{1}})$, and $M_{1}^{T}=M_{0}^{T}$. Again, $\mathcal{A}$ is schurian by Lemma 3.5.
@@
Since a section of a Schur group is a Schur group, to prove Theorem 1.4, it suffices to prove that the group
$G\cong C_{8}\times C_{2}\times C_{p},$
where $p$ is an odd prime, is not a Schur group. Let $A$, $B$, and $C$ be subgroups of $G$ such that
$A\cong C_{8},~B\cong C_{2},~P\cong C_{p},~\text{and}~G=A\times B\times P.$
Denote generators of $A$ and $B$ by $a$ and $b$, respectively. Put
$A_{1}=\langle a^{2}\rangle\cong C_{4},~A_{2}=\langle a^{4}\rangle\cong C_{2},~U_{1}=A_{1}\times P\cong C_{4p},~\text{and}~U=A_{1}\times B\times P=B\times U_{1}\cong C_{2}\times C_{4p}.$
Let us construct a nonschurian $S$-ring over $G$. The group $\operatorname{Aut}(P)$ is a cyclic group of order p1p-1. Let $M_{0}$ be the subgroup of $\operatorname{Aut}(P)$ of index $2$. The canonical epimorphism from $\operatorname{Aut}(P)$ to $\operatorname{Aut}(P)/M_{0}$ is denoted by $\pi$. Since $\operatorname{Aut}(A_{1})$ is a group of order $2$, there exists a unique isomorphism $\varphi$ from $\operatorname{Aut}(A_{1})\cong C_{2}$ to $\operatorname{Aut}(P)/M_{0}\cong C_{2}$. Put
$\mathcal{A}_{1}=\operatorname{Cyc}(M_{1},U_{1}),$
where $M_{1}=\{(\delta,\tau)\in\operatorname{Aut}(A_{1})\times\operatorname{Aut}(P):~\delta^{\varphi}=\tau^{\pi}\}\leq\operatorname{Aut}(U_{1})$, and
$\mathcal{A}_{2}=\mathcal{T}_{B}\otimes\mathcal{T}_{P}.$
@@
Put
$\mathcal{A}_{3}=\operatorname{Cyc}(\langle\delta_{0}\rangle,A),$
where $\delta_{0}\in\operatorname{Aut}(A)$ is such that $a^{\delta_{0}}=a^{-1}$. It is a straightforward to check that the partition of $(A/A_{2})\times B\cong C_{4}\times C_{2}$ into the sets
@@
defines the $S$-ring $\mathcal{A}_{4}$ over $(A/A_{2})\times B$ such that
@@
Since $X_{3}=P^{\#}$ is an orbit of $K$, we have
@@
Let $a,b,c,d\in H$ be such that $\langle a,b,c,d\rangle=H$, $A=\langle a\rangle$, $B=\langle b\rangle$, $C=\langle c\rangle$, and let $\delta_{0}\in\operatorname{Aut}(H)$ be such that
$\delta_{0}:(a,b,c,d)\mapsto(a,ab,bc,cd),$
$M_{0}=\langle\delta_{0}\rangle$, and
$\mathcal{A}_{0}=\operatorname{Cyc}(M_{0},H).$
It is straightforward to verify that $|\delta_{0}|=4$ and the basic sets of $\mathcal{A}_{0}$ are the following:
@@
One can see that each of two latter basic sets has trivial radical and generates $H$. So $\mathcal{A}_{0}$ is indecomposable. Every basic set of $\mathcal{A}_{0}$ has the $2$-power size, i.e. $\mathcal{A}_{0}$ is a 22-$S$-ring in terms of [19]. Clearly, $\mathcal{A}_{0}$ is cyclotomic and hence schurian. Therefore $\mathcal{A}_{0}$ is $2$-minimal by [19, Lemma 5.6] stating that every schurian indecomposable $p$-$S$-ring over an elementary abelian group of order $p^{4}$ is $2$-minimal for every prime $p$. Thus,
@@
for every $K_{0}\leq\operatorname{Sym}(H)$ such that $K_{0}\geq H_{r}$ and $K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0})$ (recall that $H_{r}$ is the group of $\operatorname{Sym}(H)$ induced by all right multiplications of $H$). In particular, $\operatorname{Aut}(\mathcal{A}_{0})=H_{r}\rtimes M_{0}$.
Let $V=A\times B\times C$ and $U=V\times P$. As $p\geq 5$, the group $M_{P}=\operatorname{Aut}(P)\cong C_{p-1}$ has a nontrivial subgroup $M_{P}^{0}$ of index $2$. Let $\delta_{1},\delta_{2}\in\operatorname{Aut}(V)$ be such that
$\delta_{1}:(a,b,c)\mapsto(a,ab,bc),~\delta_{2}:(a,b,c)\mapsto(a,b,bc).$
Put $M_{V}=\langle\delta_{1},\delta_{2}\rangle$. It can be verified directly that $|\delta_{1}|=4$, $|\delta_{2}|=2$, and $\delta_{2}\delta_{1}\delta_{2}=\delta_{1}^{-1}$. So $M_{V}\cong D_{8}$.
Let $M^{0}_{V}=\langle\delta_{1}^{2},\delta_{2}\delta_{1}\rangle$. Clearly, $M_{V}^{0}\cong E_{4}$ and hence $M_{V}/M_{V}^{0}\cong C_{2}$. Put
$M_{1}=\{(\delta,\tau)\in M_{V}\times M_{P}:~(\delta^{\pi_{1}})^{\psi}=\tau^{\pi_{2}}\}\leq\operatorname{Aut}(U),$
where $\pi_{1}$ and $\pi_{2}$ are the canonical epimorphisms from $M_{V}$ to $M_{V}/M_{V}^{0}$ and from $M_{P}$ to $M_{P}/M_{P}^{0}$, respectively, and $\psi$ is a unique isomorphism from $M_{V}/M_{V}^{0}\cong C_{2}$ to $M_{P}/M_{P}^{0}\cong C_{2}$, and
$\mathcal{A}_{1}=\operatorname{Cyc}(M_{1},U).$
The definition of $M_{1}$ implies that $V$ and $P$ are $\mathcal{A}_{1}$-subgroups and the basic sets of $\mathcal{A}_{1}$ are the following:
@@
To complete the proof of Theorem 1.5, it suffices to prove the proposition below.
@@
$K_{1}\geq U_{r},~K_{0}\geq(G/P)_{r},~K_{1}\approx_{2}\operatorname{Aut}(\mathcal{A}_{1}),~K_{0}\approx_{2}\operatorname{Aut}(\mathcal{A}_{0}),~\text{and}~K_{0}^{S}=K_{1}^{S}$
@@
$X=AcP_{1}\cup AbcP_{2}$
and $x\in P_{1}$. Suppose that $f$ belongs to the one-point stabilizer $M_{x}$ of $x$ in $M$, where $M=(K_{1})_{e}$ is the one-point stabilizer of $e$ in $K_{1}$. Since $K_{0}^{S}=K_{1}^{S}$, Eq. (42) implies that $f^{S}\in M_{0}^{S}$. In view of Eq. (2) and $(A\times B)c,X^{-1}\in\mathcal{S}(\mathcal{A})$, we obtain
@@
Together with $f^{S}\in M_{0}^{S}$ and the definition of $M_{0}$, this shows that $f^{S}\in\langle\delta_{0}^{2}\rangle$ and hence $b^{f}=b$. Since $H$ and $P$ are $\mathcal{A}$-subgroups, $Hx$ and $Pb$ are blocks of $\operatorname{Aut}(\mathcal{A})$. As $x^{f}=x$ and $b^{f}=b$, we conclude that $(Hx)^{f}=Hx$ and $(Pb)^{f}=Pb$. Therefore
@@
Thus, $f$ belongs to the one-point stabilizer $M_{bx}$ of $bx$ in $M$ and consequently
@@
The definition of $M_{1}$ implies that the basic set of $\mathcal{A}$ containing $x$ is $P^{\#}$ and the basic set of $\mathcal{A}$ containing $bx$ is $Y=AbP^{\#}$. Clearly, |Y|=2(p1)|Y|=2(p-1). By the assumption, we have $P^{\#},Y\in\operatorname{Orb}(M,U)$. Therefore
@@
in contrast to Eq. (43). ∎
Acknowledgment: The author is very grateful to Prof. Ilia Ponomarenko for a careful reading and valuable comments which help to improve the text significantly. The author would like to thank Dr. Matan Ziv-Av for the help with computer calculations.